# Polarized $W^{+}W^{-}$ pairs at the LHC: Effects from bottom-quark induced processes at NLO QCD+EW

### v1

@@ Abstract
We investigate the effects of the bottom-quark induced processes on the doubly polarized cross sections of $W^{+}W^{-}$ pair production at the LHC. The method to extract the on-shell single-top contribution is provided. Results for phenomenological and experimental analyses are given at next-to-leading order (NLO) QCD+EW accuracy, with the leading contribution from the gluon-gluon and photon-photon fusion included. We found that the contribution of the bottom-quark induced processes, after the subtraction of the on-shell $tW$ channel, is largest for the doubly longitudinal polarization, being at the level of $9\%$ and $13\%$ (compared to the NLO value of the light-quark contribution, for integrated cross section) for two cut setups, with and without a jet veto, respectively. A bound of the $tW$ interference is calculated for various kinematic distributions, showing that this interference effect is, in general, smaller for the no jet veto case.
@@ 1 Introduction
In parallel, active theoretical efforts to provide more precise theoretical predictions including next-to-leading order (NLO) QCD and EW corrections have been going on. Fixed-order results for doubly-polarized cross sections for $ZZ$ Denner:2021csi , $W^{\pm}Z$ Denner:2020eck ; Le:2022lrp ; Le:2022ppa , $W^{+}W^{-}$ Denner:2020bcz ; Denner:2023ehn ; Dao:2023kwc have been obtained at NLO including both QCD and EW corrections for fully leptonic decays. Next-to-next-to-leading order (NNLO) QCD results for $W^{+}W^{-}$ were provided in Poncelet:2021jmj , taking into account the light quark induced processes (without the bottom quark contribution). Semileptonic final state has been considered in Denner:2022riz for the case of $WZ$ at NLO QCD.
@@
Indeed, even for the case of unpolarized $W^{+}W^{-}$ pair production, we are not aware of any NLO theoretical work taking into account the bottom-quark induced processes with the on-shell $tW$ production subtracted. In Bierweiler:2012kw ; Baglio:2013toa , the bottom-quark induced processes are neglected assuming a perfect $b$ tagging efficiency. In Biedermann:2016guo , the $b\gamma$ processes are included, but the on-shell $tW$ contribution was not subtracted. It was pointed out in Baglio:2013toa (where one of us is an author, see Section 3.6 there) that the contribution of the bottom-quark induced processes is very small after the subtraction of the on-shell $tW$ production. This was an additional argument for neglecting the bottom-induced processes.
@@ 3 The $tW$ contribution
Figure 2: Bottom-quark induced processes at NLO QCD: virtual corrections (a), real-gluon emission (b) and bottom emission (c,d), where (d) is the $tW$ contribution. Similar diagrams occur for the NLO EW corrections, where the gluon is replaced by the photon.
@@
By now, in most ATLAS and CMS analyses of the $WW$ process, the $tW$ channel is part of the background. We will therefore consider two scenarios: with and without the $tW$ contribution. Subtracting the $tW$ will remove both the resolved and unresolved bottom quark contribution of the $tW$ origin. The signal includes the $WW$ diagrams ((a), (b), and (c) of Fig. 2) and the interference between diagrams (c) and (d) of Fig. 2.
This interference term cannot be calculated in a gauge-invariant way  [^1] A method was proposed in Ref. Frixione:2008yi to estimate the interference effect from the difference between the on-shell subtraction and diagram removal results. This is not gauge invariant because the diagram removal result is not gauge invariant. However, its bound can be estimated as follows. Since this interference effect occurs only in the $bg$ (for QCD corrections) and $b\gamma$ (EW corrections) processes, we consider only the $bg$ and $b\gamma$ processes ($b$ here can be the bottom or anti-bottom quark). We first calculate the full (including $WW$, $tW$ and their interference) result, then subtract from this the on-shell $tW$ contribution. There remains the sum of the $WW$ and the interference contributions, named $\hat{\sigma}_{\text{TW-int}}$ in the following (the hat symbol is to indicate this is a bound). Since there is no method to separate the interference term from the $WW$ one in a gauge-invariant way, we therefore interpret the absolute value of $\hat{\sigma}_{\text{TW-int}}$ as the bound of the interference effect. We will keep the sign to show that it can be negative.
Technically, using the dipole-subtraction method Catani:1996vz ; Dittmaier:1999mb , $\hat{\sigma}_{\text{TW-int}}$ includes a contribution from the so-called $(n+1)$-particle part (where the extra $b$-radiation amplitudes and subtraction terms are added) and the corresponding PK-operator term (where the integrated part and PDF counterterms are included). Note that the virtual corrections and the I-operator (or endpoint) terms (groups (a) and (b)) do not contribute to the $tW$ interference term.
@@ 4 Numerical results
We employ here the identical input parameters and renormalization schemes (for QCD and EW corrections) as in our previous investigations Le:2022ppa ; Dao:2023pkl . Note that $m_{b}=4.7{\,\text{GeV}}$ for the loop-induced gluon gluon contribution, otherwise the bottom quark is massless. The top-quark parameters are $m_{t}=173{\,\text{GeV}}$, $\Gamma_{t}=1.42{\,\text{GeV}}$. We discuss numerical results for the LHC at $13$ TeV center-of-mass energy. The factorization and renormalization scales are chosen at a fixed value $\mu_{F}=\mu_{R}=M_{W}$, where $M_{W}=80.385\,{\,\text{GeV}}$. For the parton distribution functions, we use the Hessian set LUXqed17_plus_PDF4LHC15_nnlo_30 Watt:2012tq ; Gao:2013bia ; Harland-Lang:2014zoa ; Ball:2014uwa ; Butterworth:2015oua ; Dulat:2015mca ; deFlorian:2015ujt ; Carrazza:2015aoa ; Manohar:2016nzj ; Manohar:2017eqh via the library LHAPDF6 Buckley:2014ana . For the PDF counterterms, the DIS scheme (see e.g. Dittmaier:2009cr ) has been used in the calculation of the NLO EW corrections. We have checked that the difference compared to the $\overline{\text{MS}}$ scheme (see e.g. Dittmaier:2009cr ) is negligible.
@@ 4.1 Integrated polarized cross sections
We first present results for the unpolarized and four polarized (LL, LT, TL, TT) integrated cross sections in Table 1. The polarization interference (denoted Pol-int), calculated by subtracting the polarized cross sections from the unpolarized one, is shown in the bottom row. Besides the three defined cross sections, $\sigma_{\text{NoB}}$, $\sigma_{\text{YesTW}}$ and $\sigma_{\text{NoTW}}$, with their statistical and scale uncertainties, we show some ratios. We choose to normalize new effects to the $\sigma_{\text{NoB}}$ to quantify various bottom-induced corrections. The LO $b\bar{b}$ contributions is quantified by a K-factor, $K^{\text{LO}}_{b\bar{b}}=(\sigma_{\text{NoB}}+\sigma_{b\bar{b}}^{\text{LO}})/\sigma_{\text{NoB}}$ for unpolarized and each polarized cross sections. Similarly for the NLO contribution to the b-induced processes, we give also the K-factors defined as: $K_{\text{NoTW}}=\sigma_{\text{NoTW}}/\sigma_{\text{NoB}}$; $K_{\text{YesTW}}=\sigma_{\text{YesTW}}/\sigma_{\text{NoB}}$. The interference effect is quantified by $\hat{\delta}_{\text{TW-int}}=\sigma_{\text{TW-int}}/\sigma_{\text{NoB}}$. The last three columns of Table 1, we give the corresponding polarization fractions $f_{X,i}=\sigma_{X,i}/\sigma_{X,\text{Unpol.}}$ with $X=$ NoB, NoTW, YesTW and $i=$ Unpol., LL, LT, TL, TT, Pol-int.
@@ 4.2 Kinematic distributions
Figure 3: Distributions in $\cos\theta^{VV}_{e^{+}}$ for the YesVeto (left) and NoVeto (right) setups. The angle is calculated in the $VV$ center-of-mass system (more details are provided in the text), hence denoted with the $VV$ superscript. The big panel shows the absolute values of the NoTW cross sections. The small panels, from top to bottom, are the EW $tW$-interference bound, the QCD $tW$-interference bound, the ratio of $\sigma_{\text{NoTW}}/\sigma_{\text{NoB}}$, the ratio of $\sigma_{\text{YesTW}}/\sigma_{\text{NoB}}$, the normalized shapes of the distributions shown in the big panel.
@@
The effect from b-quark induced processes on the doubly longitudinal polarization is also interesting when considering differential cross sections. In Fig. 3, we present the distributions in $\cos(\theta_{e}^{\text{WW}})$. The left plot shows the distributions with the YesVeto setup, while the right plot shows the distributions with NoVeto setup. Here, $\theta_{e}^{\text{WW}}$ is defined as the angle between the momentum vectors of the positron ($\vec{p}_{e^{+}}$) and the $W^{+}$ boson ($\vec{p}_{W^{+}}$). The positron momentum is calculated in the rest frame of the $W^{+}$ boson, while the $W^{+}$ momentum is defined in the $WW$ frame. The top panels of Fig. 3 display the absolute values of the unpolarized and polarized cross sections, denoted as $\sigma_{\text{NoTW}}$, which include all contributions except the on-shell $tW$ channel. The polarization interference effect is the difference between the unpolarized (black) and the polarization sum (pink) lines.
@@
On the next two small panels, we show the $tW$-inteference effect ($\hat{\sigma}_{\text{TW-int}}$) of the EW and QCD contributions separately. For the YesVeto case, both EW and QCD contributions are mostly negative across the whole range of $\cos(\theta_{e}^{\text{WW}})$. The QCD contribution to the LL polarization is the most negative, it however is less than $10\%$ in magnitude and distributed rather evenly over the range. The distributions are rather different for the NoVeto case. Both EW and QCD contributions can be positive and close to zero for $|\cos(\theta_{e}^{\text{WW}})|<0.75$, but grow increasingly significant at the two edges. In particular, EW and QCD contributions to TL, TT polarizations are positive while those to LT, LL are negative at the edges. There is a cancellation between the negative and positive contributions, that explains the small and positive value of the $tW$-interference term in the integrated cross sections for the NoVeto case. We also remark that the polarization interference is small for both YesVeto and NoVeto setups, and is smaller for the latter one. This interference is not always small, see e.g. the $\Delta\phi_{e,\mu}$ distribution in Fig. 7.
@@
The two following panels display the two respective K-factors ($K_{\text{NoTW}}$, $K_{\text{YesTW}}$) as function of $\cos(\theta_{e}^{\text{WW}})$. As inferred from the graphs, the K-factor for the polarization (TT) is nearly flat and close to one across the range. It is more interesting for the three other polarizations, especially the LL one. The $K_{\text{NoTW}}$ is less than one for $\cos(\theta_{e}^{\text{WW}})<-0.85$ since the $tW$-interference is more negative in this region. It is however slightly over $1.0$ for the rest. It is more striking for the YesTW plot. The K-factor is larger than two and stands out from the other polarizations. Similarly for the case of NoVeto setup, however the K-factors of the doubly longitudinal polarization are even larger than the YesVeto case. Its maximum reaches 7 around $\cos(\theta_{e}^{\text{WW}})=-0.8$ for the YesTW result.
@@
In the bottom panels of the $\cos(\theta_{e}^{\text{WW}})$ distributions, we display the normalized distributions from the top panel where the integrated cross sections are all normalized to unity, to highlight the shape differences.
@@
We will now discuss the transverse momentum distributions of the $W^{+}$ in Fig. 4. The distributions are presented similarly to the $\cos(\theta_{e}^{\text{WW}})$ distributions, with YesVeto setup on the left and NoVeto setup on the right. The $tW$-inteference effects of the EW contributions ($\hat{\sigma}_{\text{TW-int}}^{\text{EW}}$) for all polarizations are not interesting, they are close to zero and show a slight dependence on $p_{T,W^{+}}$. However $\hat{\sigma}_{\text{TW-int}}^{\text{QCD}}$ in the YesVeto case is more interesting. While the TT polarization does not depend on $p_{T,W^{+}}$, the mix polarizations (TL, LT) show a similar dependence and the LL polarization exhibits a strongest dependence. In particular, the QCD contribution to the $tW$-inteference effect for LL polarization is negative and and starting from $p_{T,W^{+}}>50\,{\,\text{GeV}}$ it falls down rapidly and reaches $-20\%$ at $p_{T,W^{+}}\approx 200\,{\,\text{GeV}}$. This behavior of $\hat{\sigma}_{\text{TW-int}}^{\text{QCD}}$ does not occur for the NoVeto setup, and the $tW$-interference is much smaller there. The smallness of this interference in the NoVeto case is also visible for all distributions displayed in Appendix B. This result suggests that the NoVeto setup seems to be a better choice for precise polarization measurements.
@@
The dependence of the two K-factors ($K_{\text{NoTW}}$, $K_{\text{YesTW}}$) for the LL polarization is also interesting. The $K_{\text{NoTW}}$ line curves down and peaks around $p_{T,W^{+}}=125$ GeV ($150$ GeV) for YesVeto (NoVeto), while, including the on-shell $tW$ contribution, $K_{\text{YesTW}}$ increases fast with $p_{T,W^{+}}$ for both cut setups.
@@
We have a technical note here. To have a smooth plot for the $tW$-interference is not easy as it is the result of a subtraction of two huge numbers whose values are very close. One can see this from the fluctuations in the distributions, in particular Fig. 4 and Fig. 8. In order to obtain the shapes as shown in this paper, we have doubled the statistics (for both individual runs and the number of random seeds) for the components contributing to the $tW$-interference. Specifically, $20$ random seeds, each with $2^{24}$ random points (for the $(n+1)$ QCD components, $2^{22}$ for the corresponding EW) have been used to calculate the $tW$-interference distributions. Usually, choosing $10$ random seeds with $2^{23}$ QCD points is enough to have nice distributions.
Finally, additional kinematic distributions, which are important for polarization separation, are provided in Appendix B.

### v2

@@ Abstract
We investigate the effects of the bottom-quark induced processes on the doubly polarized cross sections of $W^{+}W^{-}$ pair production at the LHC. The method to extract the on-shell single-top contribution is provided. Results for phenomenological and experimental analyses are given at next-to-leading order (NLO) QCD+EW accuracy, with the leading contribution from the gluon-gluon and photon-photon fusion included. We found that the contribution of the bottom-quark induced processes, after the subtraction of the on-shell $tW$ channel, is largest for the doubly longitudinal polarization. At the integrated cross section level, using a fiducial ATLAS cut with a jet veto, the effect is $9\%$ compared to the NLO value of the light-quark contribution. It increases to $13\%$ after removing the jet veto. A bound of the $tW$ interference is calculated for various kinematic distributions, showing that this interference effect is, in general, smaller for the no jet veto case. Relevant scale uncertainties are calculated to help us decide on the importance of this interference.
@@ 1 Introduction
In parallel, active theoretical efforts to provide more precise theoretical predictions including next-to-leading order (NLO) QCD and EW corrections have been going on. Fixed-order results for doubly-polarized cross sections for $ZZ$ Denner:2021csi , $W^{\pm}Z$ Denner:2020eck ; Le:2022lrp ; Le:2022ppa , $W^{+}W^{-}$ Denner:2020bcz ; Denner:2023ehn ; Dao:2023kwc have been obtained at NLO including both QCD and EW corrections for fully leptonic decays. Next-to-next-to-leading order (NNLO) QCD results for $W^{+}W^{-}$ were provided in Poncelet:2021jmj , taking into account the light quark induced processes (without the bottom quark contribution). Semileptonic final state has been considered in Denner:2022riz for the case of $WZ$ at NLO QCD. NLO QCD+EW corrections have been very recently calculated for the same-sign $W^{+}W^{+}jj$ production in Denner:2024tlu , marking the first NLO calculation of polarized vector-boson scattering.
@@
For the case of unpolarized $W^{+}W^{-}$ pair production, it was pointed out in Baglio:2013toa (where one of us is an author, see Section 3.6 there) that the contribution of the bottom-quark induced processes is very small at NLO after the subtraction of the on-shell $tW$ production. This conclusion is supported by a detailed NNLO QCD calculation Gehrmann:2014fva , showing that the $b$-induced effect (after the subtraction of the on-shell top processes) is about $2\%$ at $\sqrt{s}=14$ TeV. This was an additional argument for neglecting the bottom-induced processes.
@@
The above approach, which will be adopted in this paper, is well known in the literature as the five-flavor scheme (5FS). Alternatively, one can use the four-flavor scheme (4FS) with a massive $b$ quark to calculate the bottom-induced effects. In the 4FS, the $b$ quark is treated differently from the other light quarks in the following aspects: (i) The bottom parton distribution function (PDF) is zero, hence the $b$ quark is absent in the initial state. (ii) A finite value of $m_{b}$ serves as an infrared (IR) cutoff, allowing for an alternative method to explore the IR region. In the 4FS, the $b$-induced effects in $W^{+}W^{-}$ pair production can be calculated using the $W^{+}W^{-}b\bar{b}$ matrix elements. This final state includes off-shell $t\bar{t}$, $tW$, and $W^{+}W^{-}$ in association with $b\bar{b}$ production processes and their interference. As in the case of the 5FS, we have to subtract from this coherent sum the on-shell top contribution to obtain the “top-free” $W^{+}W^{-}$ signal. Unpolarized cross section results for the $W^{+}W^{-}b\bar{b}$ production in the 4FS can be found in Frederix:2013gra ; Cascioli:2013wga (at NLO QCD) and in Jezo:2016ujg ; Jezo:2023rht (with parton shower effects).
@@ 3 The $tW$ contribution
Figure 2: Bottom-quark induced processes at NLO QCD in the 5FS: virtual corrections (a), real-gluon emission (b) and bottom emission (c,d), where (d) is the $tW$ contribution. Similar diagrams occur for the NLO EW corrections, where the gluon is replaced by the photon.
@@
In this work, we use the 5FS, where the bottom quark is approximated as massless (except for the loop-induced gluon-gluon fusion where we take $m_{b}=4.7$ GeV) and the bottom PDF of the proton is non-vanishing. At NLO, the bottom quark induced processes include all processes with at least one $b$ quark in the initial state, as shown in Fig. 2. At NNLO, where two additional jets are possible, processes with two $b$ quarks in the final state such as $gg\to W^{+}W^{-}b\bar{b}$ and $q\bar{q}\to W^{+}W^{-}b\bar{b}$ (with $q=u,d,c,s$) are also counted as $b$-induced processes.
By now, in most ATLAS and CMS analyses of the $W^{+}W^{-}$ process, the top contribution is part of the background. Subtracting the $t$ backgrounds will remove both the resolved and unresolved bottom quark contribution of the top origin. At NLO, the signal includes the $WW$ diagrams ((a), (b), and (c) of Fig. 2) and the interference between diagrams (c) and (d) of Fig. 2.
This interference term cannot be calculated in a gauge-invariant way  [^1] A method was proposed in Ref. Frixione:2008yi to estimate the interference effect from the difference between the on-shell subtraction and diagram removal results. This is not gauge invariant because the diagram removal result is not gauge invariant.. However, its bound can be estimated as follows. Since this interference effect occurs only in the $bg$ (for QCD corrections) and $b\gamma$ (EW corrections) processes, we consider only the $bg$ and $b\gamma$ processes ($b$ here can be the bottom or anti-bottom quark). We first calculate the full (including $WW$, $tW$ and their interference) result, then subtract from this the on-shell $tW$ contribution. There remains the sum of the $WW$ and the interference contributions, named $\hat{\sigma}_{\text{TW-int}}$ in the following (the hat symbol is to indicate this is a bound). Since there is no method to separate the interference term from the $WW$ one in a gauge-invariant way, we therefore interpret the absolute value of $\hat{\sigma}_{\text{TW-int}}$ as the bound of the interference effect. We will keep the sign to show that it can be negative. The bound $\hat{\sigma}_{\text{TW-int}}$ is gauge invariant by construction.
More specifically, for the NLO QCD contribution we have
$\hat{\sigma}_{\text{TW-int}}=\sigma_{bg}-\sigma_{bg}^{tW},$ (4)
where the $b$ and $g$ PDFs are included and the second term is the on-shell $tW$ contribution detailed in Appendix A. $\sigma_{bg}$ is the cross section of the process $bg\to e^{+}\nu_{e}\mu^{-}\bar{\nu}_{\mu}b$.
The $bg\to e^{+}\nu_{e}\mu^{-}\bar{\nu}_{\mu}b$ amplitude includes the diagrams (c) and (d) of Fig. 2 and all related diagrams which form a gauge invariant group. The diagrams like Fig. 2 (c) are IR divergent because of the singular $g\to b\bar{b}$ splitting. As usual, this divergence is absorbed into the PDF as explained in Catani:1996vz for the case of hadron-hadron collision.
The cross section $\sigma_{bg}$ is calculated using the dipole-subtraction method Catani:1996vz ; Dittmaier:1999mb . It includes a contribution from the so-called $(n+1)$-particle part (where the extra $b$-radiation amplitudes and subtraction terms are combined) and the corresponding PK-operator term (where the integrated part and PDF counterterms are included). These two contributions are individually finite and hence can be separately calculated using an IR regularization, for which we choose the mass regularization Dittmaier:1999mb . Similarly, the $tW$ interference bound for the NLO EW contribution can be straightforwardly obtained.
@@ 4 Numerical results
We employ here the identical input parameters and renormalization schemes (for QCD and EW corrections) as in our previous investigations Le:2022ppa ; Dao:2023pkl . Note that $m_{b}=4.7{\,\text{GeV}}$ for the loop-induced gluon gluon contribution, otherwise the bottom quark is massless. The top-quark parameters are $m_{t}=173{\,\text{GeV}}$, $\Gamma_{t}=1.42{\,\text{GeV}}$. We discuss numerical results for the LHC at $13$ TeV center-of-mass energy. The factorization and renormalization scales are chosen at a fixed value $\mu_{F}=\mu_{R}=M_{W}$, where $M_{W}=80.385\,{\,\text{GeV}}$. For the parton distribution functions, we use the Hessian set LUXqed17_plus_PDF4LHC15_nnlo_30 Watt:2012tq ; Gao:2013bia ; Harland-Lang:2014zoa ; Ball:2014uwa ; Butterworth:2015oua ; Dulat:2015mca ; deFlorian:2015ujt ; Carrazza:2015aoa ; Manohar:2016nzj ; Manohar:2017eqh via the library LHAPDF6 Buckley:2014ana . For the PDF counterterms, the DIS scheme (see e.g. Dittmaier:2009cr ) has been used in the calculation of the NLO EW corrections. We have checked that the difference compared to the $\overline{\text{MS}}$ scheme (see e.g. Dittmaier:2009cr ) is negligible. For the NLO QCD corrections, the $\overline{\text{MS}}$ scheme is used.
@@ 4.1 Integrated polarized cross sections
We first present results for the unpolarized and four polarized (LL, LT, TL, TT) integrated cross sections in Table 1. The polarization interference (denoted Pol-int), calculated by subtracting the polarized cross sections from the unpolarized one, is shown in the bottom row. Besides the three defined cross sections, $\sigma_{\text{NoB}}$, $\sigma_{\text{YesTW}}$ and $\sigma_{\text{NoTW}}$, with their statistical and scale uncertainties, we show some ratios. We choose to normalize new effects to the $\sigma_{\text{NoB}}$ to quantify various bottom-induced corrections. The LO $b\bar{b}$ contributions is quantified by a R-factor, $R^{\text{LO}}_{b\bar{b}}=(\sigma_{\text{NoB}}+\sigma_{b\bar{b}}^{\text{LO}})/\sigma_{\text{NoB}}$ for unpolarized and each polarized cross sections. Similarly for the NLO contribution to the b-induced processes, we give also the R-factors defined as: $R_{\text{NoTW}}=\sigma_{\text{NoTW}}/\sigma_{\text{NoB}}$; $R_{\text{YesTW}}=\sigma_{\text{YesTW}}/\sigma_{\text{NoB}}$. The interference effect is quantified by $\hat{\delta}_{\text{TW-int}}=\hat{\sigma}_{\text{TW-int}}/\sigma_{\text{NoB}}$. The last three columns of Table 1, we give the corresponding polarization fractions $f_{X,i}=\sigma_{X,i}/\sigma_{X,\text{Unpol.}}$ with $X=$ NoB, NoTW, YesTW and $i=$ Unpol., LL, LT, TL, TT, Pol-int.
@@ 4.2 Kinematic distributions
Refer to caption Refer to caption
Figure 3: Distributions in $\cos\theta^{WW}_{e^{+}}$ for the YesVeto (left) and NoVeto (right) setups. Details are provided in the main text.
Refer to caption Refer to caption
@@
The effects from $b$-quark induced processes on the doubly longitudinal polarization are also interesting when considering differential cross sections. We first discuss two important distributions, an angular distribution of the electron and a transverse momentum distribution of the $W^{+}$, in Fig. 3 and Fig. 4, respectively. Other distributions Figs. 5-8, which are also important for polarization separation, are shown in Appendix B.
For each distribution we display four plots arranged in two columns and two rows. The first column is for the YesVeto setup, the second one NoVeto. Each plot on the first row has six panels. The big panel shows the values of the NoTW unpolarized and polarized cross sections, which include all contributions except the on-shell $tW$ channel. The small panels, from top to bottom, are the EW $tW$-interference bound, the QCD $tW$-interference bound, the ratio of $\sigma_{\text{NoTW}}/\sigma_{\text{NoB}}$, the ratio of $\sigma_{\text{YesTW}}/\sigma_{\text{NoB}}$, the normalized shapes of the distributions shown in the big panel. Each plot on the second row has six panels showing the scale uncertainties. From top to bottom, the first three panels show the relative scale uncertainties of the NoTW cross sections provided in the big panel on the first row. The next three panels display the scale uncertainty band of the EW+QCD $tW$ interference bound $\hat{\delta}_{\text{TW-int}}=\hat{\sigma}_{\text{TW-int}}/\sigma_{\text{NoB}}$. This band is calculated by varying the scales $\mu_{F}$ and $\mu_{R}$ in $\hat{\sigma}_{\text{TW-int}}$ while $\sigma_{\text{NoB}}$ is fixed at the central scale $\mu_{F}=\mu_{R}=M_{W}$. Same as for the integrated cross sections, the seven-scale method is used to compute the scale uncertainty band.
In Fig. 3, we present the distributions in $\cos(\theta_{e}^{\text{WW}})$. Here $\theta_{e}^{\text{WW}}$ is defined as the angle between the momentum vectors of the positron ($\vec{p}_{e^{+}}$) and of the $W^{+}$ boson ($\vec{p}_{W^{+}}$). The positron momentum is defined in the rest frame of the $W^{+}$ boson, while the $W^{+}$ momentum in the $WW$ frame.
From top to bottom, in the big panel the TT, TL, LT, LL polarized differential cross sections together with their sum (labeled polar. sum) are plotted. The unpolarized result is also shown in grey. The difference between the unpolarized and the polarization sum (pink) lines gives the interference between the TT, TL, LT and LL amplitudes. This polarization interference is small for both YesVeto and NoVeto setups, and is smaller for the latter. It is however not always small, see e.g. the $\Delta\phi_{e,\mu}$ distribution in Fig. 7.
@@
In the next two small panels, we show the $tW$-interference effect ($\hat{\sigma}_{\text{TW-int}}$) of the EW and QCD contributions separately. As expected, the QCD contribution is completely dominant over the whole range. For the YesVeto case, both EW and QCD interference bounds are mostly negative across the whole range of $\cos(\theta_{e}^{\text{WW}})$. The bound is largest for the LL polarization, ranging from $-15\%$ to $-5\%$, being most pronounced at the edges. The distributions are very different for the NoVeto case. Both EW and QCD bounds can be positive and move closer to zero for $|\cos(\theta_{e}^{\text{WW}})|<0.75$ (see the QCD LL in particular). However, at the two edges, the magnitude of the interference bounds becomes somewhat larger for the LL and LT cases. For the TT and TL polarizations, the QCD bounds are positive at the two edges. For the LL case, we see clearly a cancellation between the negative and positive contributions in different regions, explaining the small and positive value of the $tW$-interference bound for the integrated cross section of the NoVeto setup.
@@
The two following panels display the two respective R-factors ($R_{\text{NoTW}}$, $R_{\text{YesTW}}$) as function of $\cos(\theta_{e}^{\text{WW}})$, showing the bottom-induced effects in comparison to the contribution of the two first generations. As seen from the NoTW panels for the YesVeto setup, the R-factor of the TT polarization is nearly flat and close to $1$ across the range. It is more interesting for the other three polarizations, especially the LL one. The $R_{\text{NoTW}}$ is less than $1$ for $\cos(\theta_{e}^{\text{WW}})<-0.85$ since the $tW$ interference bound is more negative in this region. It is however slightly over $1$ for the rest. The picture is completely different for the YesTW plot. The LL R-factor is mostly greater than $2$ and stands out from the other polarizations. Similarly for the case of NoVeto setup, the R-factors of the doubly longitudinal polarization are even larger than the corresponding ones of the YesVeto case. Its maximum reaches $7$ around $\cos(\theta_{e}^{\text{WW}})=-0.8$ for the YesTW case.
@@
In the next panels of the $\cos(\theta_{e}^{\text{WW}})$ distribution, we display the normalized distributions from the top panel where the integrated cross sections are all normalized to unity, to highlight the shape differences. We see that the LT and LL polarizations are completely different from the TL and TT ones. This gives additional power to distinguish polarized signals.
The relative scale uncertainties of the NoTW differential cross sections plotted in the top panel are displayed in the next three panels. The remarkable feature is the difference between the YesVeto and NoVeto results, showing that the scale uncertainties of the former are significantly smaller for all polarizations. We note that the scale uncertainties in both cases are calculated in the same way, being straightforwardly obtained from the results of the seven scale points. This result is also reflected in the integrated cross sections shown in Table 1 and Table 3. Scale uncertainties for the YesVeto setup are also provided in Denner:2020bcz for the integrated cross sections using a different PDF set (the NNPDF3.1). For the NoVeto case, the reader can consult the $WZ$ results in Denner:2020eck ; Le:2022lrp , which are expected to be close to the $W^{+}W^{-}$ ones. These results are quite close to the values in Table 1 and Table 3.
The smallness of the scale uncertainties in the YesVeto case is due to a cancellation between the inclusive zero-jet cross section and the inclusive one-jet cross section, as shown in Stewart:2011cf . The magnitude of this cancellation depends on the value of $p_{T,\text{veto}}$, which is 35 GeV in this work. It turns out that the cancellation is very strong around this value (see Fig. 1 (bottom left) of Stewart:2011cf ).
An alternative method to calculate the scale uncertainties for the YesVeto case is using the quadrature sum $\sqrt{(\Delta\sigma_{N_{j}\geq 0})^{2}+(\Delta\sigma_{N_{j}\geq 1})^{2}}$ where $N_{j}$ is the number of jets, as proposed in Stewart:2011cf . The underlying assumption here is that the two scale uncertainties $\Delta\sigma_{N_{j}\geq 0}$ and $\Delta\sigma_{N_{j}\geq 1}$ are independent. The YesVeto scale uncertainties calculated in this way are always larger than the NoVeto ones. Care must therefore be taken when using the YesVeto scale uncertainties presented in this work for comparison with data, as the values may be underestimated. This is a well-known issue of the jet-veto method.
The last three panels show the scale uncertainties of the QCD+EW $tW$-interference bound. The interesting feature here is that the scale uncertainty bands cover the line of vanishing interference for all distributions (including the ones in Appendix B). The uncertainty-band width increases with the magnitude of the interference bound. Noticeable exception is the edges of the $\cos(\theta_{e}^{\text{WW}})$ distribution for the LL case. However, the scale uncertainties of the signal are large in these regions (see the $\Delta_{\text{NoTW}}$ panel). This result shows that the $tW$-interference is, at the current level of precision, consistent with zero within the scale uncertainties. Note that, the scale uncertainties of the $tW$-interference bound are calculated here at LO. The importance of this interference effect will become clearer when the scale uncertainties are significantly reduced by including higher-order corrections.
@@
We now discuss the transverse momentum distributions of the $W^{+}$ in Fig. 4. The distributions are presented similarly to the $\cos(\theta_{e}^{\text{WW}})$ distributions, with YesVeto setup on the left and NoVeto setup on the right. The $tW$-inteference effects of the EW contributions ($\hat{\sigma}_{\text{TW-int}}^{\text{EW}}$) for all polarizations are not interesting, they are close to zero and show a slight dependence on $p_{T,W^{+}}$. However $\hat{\sigma}_{\text{TW-int}}^{\text{QCD}}$ in the YesVeto case is more interesting. While the TT polarization does not depend on $p_{T,W^{+}}$, the mixed polarizations (TL, LT) show a similar dependence and the LL polarization exhibits a strongest dependence. In particular, the QCD contribution to the $tW$-inteference bound for the LL polarization is negative and, starting from $p_{T,W^{+}}>50\,{\,\text{GeV}}$, falls down rapidly before reaching $-20\%$ at $p_{T,W^{+}}\approx 200\,{\,\text{GeV}}$. This behavior of $\hat{\sigma}_{\text{TW-int}}^{\text{QCD}}$ does not occur for the NoVeto setup. The $tW$-interference bound is much smaller there. The smallness of this interference in the NoVeto case is also visible for all distributions displayed in Appendix B. This result suggests that the NoVeto setup seems to be a better choice for precise polarization measurements.
@@
The dependence of the two R-factors ($R_{\text{NoTW}}$, $R_{\text{YesTW}}$) for the LL polarization is also interesting. The $R_{\text{NoTW}}$ line curves down and peaks around $p_{T,W^{+}}=125$ GeV ($150$ GeV) for YesVeto (NoVeto), while the $R_{\text{YesTW}}$ (including the on-shell $tW$ contribution) increases fast with $p_{T,W^{+}}$ for both cut setups.
@@
Finally, we have a technical note here. To have a smooth plot for the $tW$-interference is not easy as it is the result of a subtraction of two huge numbers whose values are very close. One can see this from the fluctuations in the distributions, in particular Fig. 4 and Fig. 8. In order to obtain the shapes as shown in this paper, we have doubled the statistics (for both individual runs and the number of random seeds) for the components contributing to the $tW$-interference. Specifically, $20$ random seeds, each with $2^{24}$ random points (for the $(n+1)$ QCD components, $2^{22}$ for the corresponding EW) have been used to calculate the $tW$-interference distributions. Usually, choosing $10$ random seeds with $2^{23}$ QCD points is enough to have nice distributions.